Question: A study of racial differences of breast cancer diagnosis was done

White Black Total
Local 45 13 58
Regional 76 9 85
Distant 35 7 42
total 156 29 185

1a)What is the conditional probability that the disease stage of a randomly selected patient is Regional, given that the patient is black?

1b)What is the conditional probability that a randomly selected patient is white, given that the patient has the disease stage Distant ?

1c)If you want to test within white breast cancer patients, are the disease stages equally distributed across “local”, “regional”, and “Distant” stages? What is the appropriate test to use:

- Two correlated sample- T test

- Two sample independent-T test

- X2 test of independence

- X2 goodness of fit test

1d)  If you want to test the disease stage and race are independent. Which of the following is the appropriate Test and test statistics value?

- X2 test of independence: X2 = 7.554

- X2 test of independence: X2 = 3.687   

- X2 goodness of fit test: X2 = 7.554

- X2 goodness of fit test: X2 = 3.687

1e) What is the P-value and df?
Answer #1

a) conditional probability that the disease stage of a randomly selected patient is Regional, given that the patient is black

=9/29=0.3103

b)conditional probability that a randomly selected patient is white, given that the patient has the disease stage Distant =35/42=0.8333

c)

-- X2 goodness of fit test

d)

Applying chi square test of independence:
Expected Ei=row total*column total/grand total local regional distant Total
white 48.9100 71.6800 35.4200 156.01
black 9.0900 13.3200 6.5800 28.99
total 58.00 85.00 42.00 185.00
chi square    χ2 =(Oi-Ei)2/Ei local regional distant Total
white 0.313 0.260 0.005 0.5779
black 1.682 1.4011 0.0268 3.1097
total 1.9944 1.6614 0.0318 3.688
test statistic X2 = 3.687

- X2 test of independence: X2 = 3.687    

e)

degree of freedom(df) =(rows-1)*(columns-1)= 2
p value = 0.1582